when c'(Q)=e^0.2Q and fixed cost is 300 find total cost,t


1
Expert's answer
2021-02-06T10:07:57-0500

Let's write the TC function:


"TC=FC+VC=300+\int C(Q)dQ,""TC=300+\int e^{0.2Q}dQ=300+5e^{0.2Q}."